Please show all work.

11. Find the transition matrix from B to B. {(1,0), (0,1)}, B {(1,1), (5,6)} B =
Answer #1

(1,0) = a[1,12 + B15, 6) =catsb, X+6B). A at SB = 1 and 2+6B=0 » B=1, 2=6 So, (1,0) = 6 11,10 + (-1)(5,6) ...(i) 10,1)= B(1,1Hence, from (i) and (is the transition materin cs Scanned with CamScanner